# Evolutionary Policy Optimization

### v1

@@ 5 Experiments
We evaluate our method on four manipulation tasks, comparing it against state-of-the-art (SOTA) approaches. Additionally, we conduct ablation studies to assess the contribution of each component. To demonstrate scalability, we analyze our method’s performance with increasing computational resources. Finally, we investigate the impact of different crossover strategies. Each aspect is discussed in detail.
@@ 5.1 Experimental Setup
Figure 2: We evaluate our algorithm on four challenging dexterous manipulation tasks from the Isaac Gym Allegro-Kuka environments, including single-arm regrasping, single-arm reorientation, two-arm regrasping, and two-arm reorientation.
Tasks
We evaluate our algorithm on four challenging dexterous manipulation tasks from the Isaac Gym Allegro-Kuka environments (Makoviychuk et al., ), including single-arm regrasping, single-arm reorientation, two-arm regrasping, and two-arm reorientation. These environments allow for diverse task completion strategies, such as throwing the object within a large exploration space, making them well-suited for evaluating the effectiveness of reinforcement learning algorithms. A visual illustration of the tasks is provided in Figure  2.
Figure 3: Performance curves of EPO with respect to PPO, PBT, PQL and SAPG baselines. In the experiment, EPO has 64 agents, and SAPG and PBT use the same number of agents as specified in their respective papers, namely 6 agents and 8 agents.
Evaluation
Following previous works (Petrenko et al., 2023; Singla et al., 2024) we use the number of successes in a single episode as a performance metric on these tasks.
  • Single-Arm Regrasping: The agent must lift a cube from the table and hold it near a randomly generated target point (shown as a white sphere in Figure 2) within 30 steps. A success is defined as the center distance between the cube and the target point being within a predefined threshold.
  • Single-Arm Throwing: The agent need pick up the cube and throw it to a bucket at a random position. A succcess is defined as if the cube is throwed into the bucket.
  • Single-Arm Reorientation: The agent must pick up an object and reorient it to a specified 6-DoF pose including both position and orientation (shown as a colored cube in Figure 2). A success is defined as both the positional and orientation differences between the object and the target pose being within a threshold.
  • Two-Arm Regrasping: The goal is the same as in one-arm regrasping. However, the generated goal pose is far from the initial pose, requiring two Kuka arms to collaborate to transfer (e.g. throwing) the object to reach target poses in different regions.
  • Two-Arm Reorientation: The goal is the same as in one-arm reorientation. However, the target pose is placed far from the initial pose, necessitating coordinated actions between the two Kuka arms.
  • Two-Arm Reorientation (Partition): Same goal as two-arm reorientation, but there is a movable partition on the table. So, the agent should first learn to remove the partition and try to put the cube in the right position.
  • Two-Arm Reorientation (Hole): Same goal as two-arm reorientation, but the barrier on the table has a hole in the middel. The agent should learn to throw the cube through the hole to successfully align the cube with its target.
  • Two-Arm Reorientation (Barrier): Same goal as two-arm reorientation, but this time the barrier on the table cannot be passed through or removed. So one allegro hand must pass the cube from the side of the barrier to another hand.
Baselines
We evaluate our approach against state-of-the-art RL methods specifically designed for the large-scale setting. Our comparisons include off-policy (Li et al., 2023) and hybrid-policy (Singla et al., 2024) variants, population-based EvoRL (Petrenko et al., 2023), as well as the standard PPO (Schulman et al., 2017).
  • Parallel Q-Learning (Li et al., 2023): A parallelized version of DDPG (Lillicrap, 2015), incorporating mixed exploration by varying exploration noise across environments to enhance exploration efficiency. This baseline is used to evaluate the scalability of off-policy algorithms as the number of samples increases. We use the official code for training without any algorithm-level modifications.
  • SAPG (Singla et al., 2024): A hybrid-policy RL algorithm designed for large-scale environments by splitting training data into chunks and recombining them via importance sampling. This baseline is used to assess the scalability of on-policy algorithms as batch size increases. We use the official code for training without any algorithm-level modifications.
  • DexPBT (Petrenko et al., 2023): A population-based training (PBT) framework built on PPO. Similarly, the $N$ environments are divided into $K$ groups, each containing $N/K$ environments, where $K$ separate agents are trained with different hyperparameters. At regular intervals, the worst-performing policies are replaced. This baseline is used to evaluate the scalability of EvoRL. We use the official code for training without any algorithm-level modifications.
  • PPO (Schulman et al., 2017): One of the most widely used reinforcement learning algorithms, serving as a strong baseline. We use this to demonstrate the scalability of standard RL.
For a fair comparison, all algorithms are trained using the same number of environments ($N=24,576$) and the same network architecture. To maintain consistency with the original papers, PBT is trained with 8 agents, and SAPG is trained with 6 agents, while EPO (ours) is trained with 64 agents. Additional ablation studies on the impact of the number of agents can be found in Section 5.3.
For simpler tasks such as regrasping, all algorithms are trained with $5\times 10^{9}$ transitions, while for more challenging tasks, all algorithms are trained with $1\times 10^{10}$ or $1.5\times 10^{10}$ transitions. Each experiment is run with 5 different random seeds, and we report the mean and standard error in the plots.
Task SAPG(Singla et al., 2024) PPO (Schulman et al., 2017) PBT (Petrenko et al., 2023) PQL (Li et al., 2023) EPO (Ours)
Single Arm Regrasping $31.6\pm 2.07$ $0.1\pm 0.11$ $24.8\pm 1.17$ $3.1\pm 0.31$ 32.1$\pm$ 1.21
Single Arm Throwing $11.5\pm 4.15$ $12.4\pm 3.11$ $9.4\pm 2.73$ $2.5\pm 0.43$ 16.4 $\pm$ 0.99
Single Arm Reorientation $26.2\pm 4.11$ $1.2\pm 0.58$ $1.9\pm 0.19$ $1.1\pm 0.47$ 36.7 $\pm$ 1.81
Two Arm Regrasping $4.1\pm 2.04$ $0.3\pm 0.15$ $6.7\pm 1.43$ $1.7\pm 0.30$ 24.1 $\pm$ 2.55
Two Arm Reorientation $2.2\pm 1.17$ $0.7\pm 0.45$ $1.7\pm 1.46$ $0.5\pm 0.22$ 35.8 $\pm$ 1.64
Two Arm Reorientation (Partition) $1.3\pm 0.94$ $0.5\pm 0.39$-- - --- - 31.8 $\pm$ 6.04
Two Arm Reorientation (Hole) $2.4\pm 0.52$ $0.5\pm 0.43$-- - --- - 24.7 $\pm$ 7.06
Two Arm Reorientation (Barrier) $1.9\pm 0.85$ $0.4\pm 0.18$-- - --- - 7.7 $\pm$ 6.5
Table 1: Mean episode success after $101e$ 10 samples ($95e$ 9 for the single-arm regrasping task) of different methods. In all tasks, our method outperforms the all baseline methods with relatively small standard deviations.
@@ 5.2 Results and Analysis
As shown in Figure 3 and Table 1, our algorithm consistently and significantly outperforms all baselines. For more visualization, please refer to https://sites.google.com/view/epo-rl.
For simpler tasks such as Kuka regrasping, SAPG and PBT perform comparably to our method, as the exploration space is relatively small. However, for more challenging tasks like single-arm reorientation, SAPG demonstrates some learning capability, whereas all other baselines fail. This highlights the advantage of incorporating genetic algorithms (GA), allowing our method to leverage large, high-quality populations to generate diverse and useful training data. Additionally, combining GA with the hybrid training strategy enables higher asymptotic performance compared to baseline methods.
For the two-arm tasks, where two Kuka robots must collaborate to complete the task, all baseline methods fail to learn any meaningful behavior even after $1\times 10^{10}$ steps, whereas EPO achieves substantial learning progress. In the original SAPG paper (Singla et al., 2024), the policy was trained for $5\times 10^{10}$ steps and converged at $28.58$ episode successes, while EPO reaches $35.8$ episode successes with only $1\times 10^{10}$ steps. This highlights the superior sample efficiency of EPO, significantly outperforming baseline methods. For the last three tasks, PPO and PQL are unable to learn any useful behaviors, so we omit their results in the figure and the result table.
Notably, EPO exhibits lower variance across different random seeds compared to baseline methods, with an average variance of $20.24\%$ of the mean rewards, while the best-performing baseline reaches $37.50\%$. This indicates that EPO is highly robust to hyperparameter variations, such as random seed initialization. This is particularly important given that reinforcement learning algorithms are notoriously sensitive to hyperparameter choices. The robustness of EPO makes it a promising approach for training RL models with minimal hyperparameter tuning, addressing a key challenge in real-world applications.
@@ 5.3 Ablations
Figure 4: Ablation of agent number ($K=8,16,32,64$) with the same total number of environments ($N=24576)$. As the number of agents increases, EPO’s performance continuously improves.
For ablation, we want to see the influence of number of population and GA. We don’t do ablation of hybrid policy update as it would be very similar as pure PBT as shown in section 5.2.
Specifically, we fix the number of environments at $N=24,576$ while increasing the population size from $K=8$ to $K=64$ (Figure 4). The results show that our algorithm benefits from a larger population. While this may seem intuitive, it is not guaranteed by default—as the number of agents increases, behavior diversity also increases, which can lead to more poorly performing agents generating noisy data. By applying Genetic Algorithms (GA) at the latent space level, EPO periodically eliminates underperforming agents, ensuring that only high-quality data is generated for the master agent to learn from. Additionally, since the actor-critic network is shared across all agents, we introduce controlled diversity while preventing unbounded divergence, which is crucial for scaling.
To further illustrate this, we remove the GA component from EPO, resulting in EPO (w/o GA), which collapses to SAPG. As shown in Figure 5, for harder tasks (e.g., two-arm manipulation), SAPG does not benefit from an increased population at all. For simpler tasks, such as single-arm regrasping, its performance even degrades compared to Figure 3. This empirically demonstrates the necessity of combining GA with hybrid-policy updates to enable effective training at scale with large datasets.
Figure 5: Comparison of the EPO w/ and w/o GA. GA helps EPO maintain the diversity of sampled data while eliminating poorly performing agents, ensuring high-quality data and accelerating training speed.
@@ 5.4 Scaling Law
Figure 6: The training curve and scaling curve (converge points)of different numbers of env. With a fixed number of environments per agent ($384$), as the total number of environments increases, EPO’s performance continues to improve.
Here, we examine the scalability of EPO with increased computational resources. Specifically, we evaluate EPO’s scalability with larger training batches generated by parallel simulations by fixing the number of environments per agent and increasing the number of agents from $K=16$ to $K=128$, thereby scaling the total number of environments from $N=6,144$ to $N=49,152$. It is worth noting that even at $K=16$, the batch size of $N=6,144$ is already significantly large. As shown in Figure 6, EPO continues to improve with increasing training data, even when the number of environments reaches $N=49,152$, demonstrating that our algorithm overcomes the limitations of existing on-policy RL methods, which typically saturate beyond a certain batch size. This highlights EPO’s strong potential for data-rich environments, such as simulations and game engines, where large training batches can be easily generated and leveraged effectively.
We also evaluate EPO’s scalability with larger neural networks, where we fix the number of agents and total environments but increase the shared network size. Specifically, we scale the network size by a factor of 4, yet observe no significant increase in per-step training time, while memory usage remains unchanged (because the network used is relatively small). This is not feasible for classic EvoRL methods, where each agent maintains its own separate weights, preventing efficient memory sharing. These results highlight EPO’s strong potential for training large-scale neural networks, including applications in LLMs (Guo et al., 2025).
@@ 6 Conclusion
Acknowledgements
We would like to thank Ananye Agarwal, Jayesh Singla, Jared Mejia, Kevin Gmelin for fruitful discussions.
Impact Statement
This paper presents Evolutionary Policy Optimization (EPO), a novel reinforcement learning algorithm designed to enhance the scalability and efficiency of policy gradient methods by integrating evolutionary search with on-policy updates. Our work primarily aims to advance the field of Machine Learning (ML) by improving the ability of reinforcement learning (RL) methods to scale with increasing computational resources, making them more effective in data-rich environments such as robotics, simulations, and game engines.
While EPO does not present direct ethical concerns, RL methods, particularly those that improve large-scale training, could have broader societal implications. Scalable RL has potential applications in autonomous systems, decision-making AI, and robotics, which may impact labor markets, automation policies, and safety regulations. It is crucial to ensure that future applications of such methods align with ethical AI principles, particularly regarding fairness, transparency, and robustness in high-stakes decision-making.
Additionally, as RL is increasingly applied to large-scale language models (LLMs) and automated reasoning systems, scalability improvements may accelerate advancements in AI alignment and human-AI interaction. However, misuse of reinforcement learning in generating persuasive AI or manipulative automated systems remains a potential risk. We encourage the community to apply EPO responsibly and consider ethical safeguards when deploying large-scale RL systems in real-world applications.

### v2

@@ 5 Experiments
Figure 2: We evaluate our algorithm on seven challenging environments that span a diverse set of tasks, including manipulation [34], locomotion [4], and classic control benchmarks [45]
We evaluate our method on eight challenging and realistic environments, including manipulation [34], locomotion [16, 4], and classic control benchmarks [45], comparing it against state-of-the-art (SOTA) approaches. Additionally, we conduct ablation studies to assess the contribution of each component. To demonstrate scalability, we analyze our method’s performance with increasing computational resources. Finally, we evaluated computational cost versus performance gain as EPO scales with agent numbers (Supplementary Material). Each aspect is discussed in detail.
@@ 5.1 Experimental Setup
Tasks We evaluate our algorithm across eight challenging environments covering a diverse range of tasks, including manipulation [34]—Single-Arm Regrasping (M:SG), Single-Arm Throwing (M:ST), Two-Arm Regrasping (M:TG), Two-Arm Reorientation (M:TR), and Two-Arm Reorientation with Obstacle (M:TO); locomotion—ANYmal Walking (L:A)[16] and Unitree A1 Parkour (L:U)[4]; and classic control—Humanoid from the DeepMind Control Suite (D)[45]. A visual overview of these tasks is shown in Figure2; additional details are provided in the Appendix.
Following prior works [34, 41], we use the number of successes per episode as the performance metric for manipulation tasks. For L:A and D, we report the cumulative rewards as defined in their original benchmarks. For L:U, we use the average terrain difficulty level achieved by all A1 agents during an episode, which serves as a proxy for success rate.
Figure 3: Performance curves of EPO compared to SAC, PQL, PPO, SAPG, PBT, and CEM-RL baselines. EPO demonstrates superior sample efficiency and higher asymptotic performance, particularly on challenging tasks.
Baselines We evaluate our approach against state-of-the-art RL methods tailored for large-scale training. The comparisons include off-policy algorithms—SAC [14] and Parallel Q-Learning [24]; on-policy methods—PPO [38]; hybrid-policy methods—SAPG [41]; and population-based EvoRL methods—PBT [34] and CEM-RL [35]. We use the original implementations of each method without introducing algorithm-level modifications. Please refer to the supplementary material for additional details.
For a fair comparison, all algorithms are trained using the same number of environments ($N=24{,}576$) and identical network architectures. For methods incorporating a population-based component (SAPG, PBT, CEM-RL), we standardize the population size to 64. Since the impact of the number of agents on final performance can vary, we also report results using each method’s original agent configuration for completeness (e.g., PBT: 8 agents; SAPG: 6 agents). Each experiment is run with 5 different random seeds, and we report the mean and standard error in the plots. Further ablation studies on the effect of population size are presented in Section 5.3.
Task SAC [14] PQL [24] PPO [38] SAPG[41] SAPG 6[41] PBT [34] PBT 8 [34] CEM-RL[41] EPO (Ours) M:SG $0.1\pm 0.00$ $3.1\pm 0.31$ $0.1\pm 0.11$ $22.4\pm 7.60$ $31.6\pm 2.07$ $6.8\pm 5.01$ $24.8\pm 1.17$ $1.7\pm 0.95$ 32.1$\pm$ 1.21 M:ST $0.0$ $2.5\pm 0.43$ $12.4\pm 3.11$ $10.8\pm 1.70$ $11.5\pm 4.15$ $2.2\pm 1.54$ $9.4\pm 2.73$ $0.0$ 16.4 $\pm$ 0.99 M:TG $0.0$ $1.7\pm 0.30$ $0.3\pm 0.15$ $1.4\pm 1.04$ $4.1\pm 2.04$ $3.1\pm 1.24$ $6.7\pm 1.43$ $0.0$ 24.1 $\pm$ 2.55 M:TR $0.0$ $0.5\pm 0.22$ $0.7\pm 0.45$ $4.4\pm 4.17$ $2.2\pm 1.17$ 0.0 $1.7\pm 1.46$ $0.0$ 35.8 $\pm$ 1.64 M:TO $0.0$ $0.0$ $0.5\pm 0.43$ $2.6\pm 0.76$ $2.4\pm 0.52$ $0.0$ $0.0$ $0.0$ 24.7 $\pm$ 7.06 L:A $64.2\pm 6.60$ $65.12\pm 3.2$ $67.7\pm 1.71$ $69.8\pm 1.81$ $68.9\pm 1.22$ $55.2\pm 1.72$ $64.7\pm 6.16$ $67.3\pm 2.01$ 74.0 $\pm$ 6.65 L:U $0.0$ $0.1\pm 0.03$ $0.6\pm 0.32$ $0.6\pm 0.45$ $0.5\pm 0.43$ $0.0$ $0.3\pm 0.31$ $0.1\pm 0.09$ 1.0 $\pm$ 0.45 D $5254.8\pm 133.95$ $8554.8\pm 593.32$ $9231.4\pm 88.47$ $11411.9\pm 790.88$ $11524.1\pm 540.23$ $7510.2\pm 117.86$ $9490.52\pm 41.09$ $7040.5\pm 103.09$ 11686.5 $\pm$ 773.7
Table 1: Performance of EPO compared to SAC, PQL, PPO, SAPG, PBT, and CEM-RL across a diverse set of tasks. In manipulation tasks, PPO, SAPG, and PBT demonstrate learning capabilities in simple scenarios; however, only EPO achieves significant learning progress on complex tasks. For locomotion tasks, where PPO is widely regarded as a strong baseline and the default method, EPO still achieves superior performance by a large margin on the more challenging problems.
Figure 4: Ablation and scaling results for EPO. (a) Ablation of agent number ($K=8,16,32,64$) with a fixed total number of environments ($N=24{,}576$). As the number of agents increases, EPO’s performance improves. (b) Training and scaling curves with increasing environments. With a fixed number of environments per agent (384), total performance improves as environments scale.
@@ 5.2 Results and Analysis
As shown in Figure 3 and Table 1, our algorithm consistently and significantly outperforms all baselines. For more visualization, please refer to https://yifansu1301.github.io/EPO/.
In manipulation tasks, while methods like PPO, SAPG, and PBT perform reasonably on simpler tasks (e.g., single-arm regrasping and throwing), EPO still achieves higher success rates. For more complex tasks requiring two-arm coordination, baselines fail to learn meaningful behavior even after $1\times 10^{10}$ steps, whereas EPO achieves strong performance. For example, EPO reaches 35.8 episode successes in two-arm reorientation using $1\times 10^{10}$ steps, surpassing SAPG’s 28.58 after $5\times 10^{10}$ steps in the original SAPG paper [41], highlighting EPO’s superior sample efficiency. A similar trend holds in locomotion. On flat-ground ANYmal walking, baselines match EPO, but in more complex parkour tasks requiring rich exploration, EPO significantly outperforms them. In the DeepMind Control Suite (Humanoid), all methods perform reasonably, yet EPO still yields marginally higher returns.
Comparing SAPG (5th column) with SAPG 6 (6th column) reveals that hybrid-policy-only updates do not benefit from large-scale simulation and may even degrade performance, as increasing the number of agents introduces a higher proportion of poorly performing policies. In contrast, comparing SAPG (5th column) with EPO (last column) highlights the critical role of the genetic algorithm (Table 1), especially in more challenging tasks like two-arm manipulation. By filtering out underperforming agents and learning only from high-quality data, the GA ensures stable and effective training. This demonstrates that the genetic algorithm is not a minor enhancement, but a fundamental component of the hybrid-policy update.
Notably, EPO demonstrates a lower performance variance across different random seeds compared to baseline methods, indicating strong robustness to hyperparameter variations, including random seed initialization. This is especially significant given the well-known sensitivity of reinforcement learning algorithms to hyperparameter settings. EPO’s robustness makes it a compelling choice for training RL models with minimal hyperparameter tuning, addressing a major challenge in deploying RL in real-world applications.
Compared to EvoRL baselines, EPO demonstrates superior learning efficiency and asymptotic performance. While PBT performs well on simpler tasks, its effectiveness diminishes in more complex scenarios. This is largely due to PBT’s requirement to train each policy nearly to convergence before selecting optimal hyperparameters, along with its inability to leverage data from past agents—resulting in inefficient use of experience. Consequently, PBT does not benefit from larger population sizes, as evidenced by the similar performance between PBT and PBT 8. In contrast, CEM-RL relies heavily on off-policy updates, which becomes a limitation in difficult tasks where most agents perform poorly. This leads to the accumulation of low-quality data and ultimately degrades asymptotic performance across a wide range of tasks. EPO overcomes these limitations by employing a hybrid-policy update mechanism while retaining the genetic algorithm for agent selection, particularly in complex environments.
@@ 5.3 Ablations
In this ablation study, we investigate the impact of population size while keeping the total computational budget fixed, using the single-arm reorientation task—where the agent must grasp an object and reorient it to a target 6-DoF pose—as the testbed. Unless otherwise specified, all subsequent analyses in this section are conducted within this environment. Specifically, we fix the number of environments to $N=24{,}576$ and vary the population size from $K=8$ to $K=64$ (Figure 4(a)). The results indicate that our algorithm benefits from a larger population. While this may seem intuitive, it is not a guaranteed outcome—larger populations introduce greater behavioral diversity, which can also increase the prevalence of poorly performing agents and result in noisier data. To address this, EPO applies Genetic Algorithms (GA) at the latent space level to periodically eliminate underperforming agents, ensuring that only high-quality data contributes to learning. Moreover, since all agents share a common actor-critic network, we maintain controlled diversity while avoiding unbounded policy divergence—an essential property for stable and scalable training.
We also analyzed the impact of different crossover strategies beyond simple averaging. Specifically, we experimented with two alternative approaches—uniform crossover and fitness-weighted averaging. In uniform crossover, each element of the latent vector is randomly selected from one of the two parents with equal probability. In the fitness-weighted average approach, the parents’ latent vectors are combined proportionally to their fitness scores. Our method achieved a 36.7% success rate, compared to 32.6% for uniform crossover and 37.0% for fitness-weighted averaging, indicating that performance is relatively robust to the choice of crossover strategy. And for simplicity, we keep the averaging strategy.
@@ 5.4 Scaling Law
Here, we examine the scalability of EPO with increased computational resources. Specifically, we evaluate EPO’s scalability with larger training batches generated by parallel simulations by fixing the number of environments per agent and increasing the number of agents from $K=16$ to $K=128$, thereby scaling the total number of environments from $N=6,144$ to $N=49,152$. It is worth noting that even at $K=16$, the batch size of $N=6,144$ is already significantly large. As shown in Figure 4(b), EPO continues to improve with increasing training data, even when the number of environments reaches $N=49,152$, demonstrating that our algorithm overcomes the limitations of existing on-policy RL methods, which typically saturate beyond a certain batch size. This highlights EPO’s strong potential for data-rich environments, such as simulations and game engines, where large training batches can be easily generated and leveraged effectively.
It is worth noting that our method focuses on scaling with training data, which is orthogonal to efforts that scale performance through larger neural network architectures. This makes it naturally compatible with existing regularization techniques designed for training large networks [32, 23]. Additionally, the shared actor-critic architecture enhances memory efficiency and inherently supports scaling to larger network sizes.
@@ 6 Conclusion
Appendix A Computational Cost
We also evaluate the computational cost of EPO compared with other baselines that involve a population-based component (SAPG, PBT, CEM-RL). Since EPO requires calculating fitness score of all agents and determining whether iteration is necessary, it slightly increases the computational cost over PBT and SAPG. However, simply increasing the number of agents in EPO does not have a significant impact on memory usage or inference time. This is because the additional cost mainly comes from adding a small number of latent embeddings, which does not drastically affect the overall resource usage. All results are obtained on task M:SO with 24,576 environments and $95e$ 9 environment steps.
Method Speed (fps) GPU Memory Usage (GB) Performance EPO (64 agents) 6764 $\sim 24$ GB 29.8 PBT (64 agents) 8120 $\sim 120$GB 1.9 SAPG (64 agents) 6672 $\sim 23$GB 13.4 CEM-RL (64 agents) 9230 $\sim 20$GB 0.0 EPO (8 agents) 6432 $\sim$21.4 GB 8.4 EPO (16 agents) 6650 $\sim$21.9 GB 21.2 EPO (32 agents) 6685 $\sim$22.9 GB 28.1 EPO (64 agents) 6764 $\sim$24GB 29.8
Appendix B Experiment Details
In this section, we present more details of our experiments.
B.1 Task and Evaluation Details
  • Manipulation: Single-Arm Regrasping (M:SG): The agent must lift a cube from the table and hold it near a randomly generated target point (shown as a white sphere in Figure 2) within 30 steps. A success is defined as the center distance between the cube and the target point being within a predefined threshold.
  • Manipulation: Single-Arm Throwing (M:ST): The agent need pick up the cube and throw it to a bucket at a random position. A succcess is defined as if the cube is throwed into the bucket.
  • Manipulation: Single-Arm Reorientation (M:SO - used for ablations): The agent must pick up an object and reorient it to a specified 6-DoF pose including both position and orientation. A success is defined as both the positional and orientation differences between the object and the target pose being within a threshold.
  • Manipulation: Two-Arm Regrasping (M:TG): The goal is the same as in one-arm regrasping. However, the generated goal pose is far from the initial pose, requiring two Kuka arms to collaborate to transfer (e.g. throwing) the object to reach target poses in different regions.
  • Manipulation: Two-Arm Reorientation (M:TR): The goal is the same as in one-arm reorientation. However, the target pose is placed far from the initial pose, necessitating coordinated actions between the two Kuka arms.
  • Manipulation: Two-Arm Reorientation with Obstacle (M:TO): The task shares the same objective as two-arm reorientation, but with an obstacle in the center of the table. The agent must learn to pass the cube through the obstacle in order to successfully align it with the target orientation.
  • Locomotion: ANYmal (L:A): A quadruped robot is tasked with following velocity commands across flat terrain.
  • Locomotion: Unitree A1 Parkour (L:U): In this task, the robot dog needs to learn how to traverse complex terrains, such as hurdles, gaps, rough terrain, etc. The robot dog will advance to the next, more difficult terrain only after it can successfully navigate the current one. Therefore, we consider the average terrain difficulty level that the robot dogs are on as the success rate. A value of 1 means that, on average, all robot dogs can complete the terrain at the first difficulty level.
  • DeepMind Control Suite: Humanoid (D): A 21-joint humanoid must run forward at a target velocity of 10 m/s without falling.
B.2 Hyperparameters and Network
Manipulation Tasks
We employ a Gaussian policy in which the mean is generated by a single-layer LSTM with 768 hidden units. The input observation is first processed by an MLP with hidden layers of sizes 768, 512, and 256, using ELU activation, before being fed into the LSTM.
Hyperparameter Value
Discount factor $0.99$
Learning rate $41e-$ 4
KL threshold for LR update $0.016$
Num Envs $24,576$
Mini-batch size 44$\cdot$4 ⋅ num_envs
Grad norm 1.0
Clipping factor 0.1
Critic coefficient 4.0
Horizon length 16
LSTM Sequence length 16
Bounds loss coefficient 0.00001
Mini Epochs 2
$\gamma$ in Eq.(3) 0.3
$\lambda$ in Eq.(5) 1.0
$x$ in Line 13 of Algo.1 num_agents$(K)-2$
Locomotion Task
We adopt a Gaussian policy in which the mean is produced by a MLP with hidden layer sizes of 512, 256, and 128, using ELU activation functions.
Hyperparameter Value
Discount factor $0.99$
Learning rate $41e-$ 4
KL threshold for LR update $0.016$
Num Envs $24,576$
Mini-batch size 44$\cdot$4 ⋅ num_envs
Grad norm 1.0
Clipping factor 0.1
Critic coefficient 4.0
Horizon length 16
Bounds loss coefficient 0.00001
Mini Epochs 2
$\gamma$ in Eq.(3) 0.5
$\lambda$ in Eq.(5) 1.0
$x$ in Line 13 of Algo.1 num_agents$(K)-2$